\begin{thebibliography}{99}

\bibitem{bb} F. J. MacWilliams and N. J. A. Sloane, The Theory of Error-Correcting
Codes. Amsterdam: North-Holland, 1977.

\bibitem{ted} Ted Hurley, MDS codes over finite fields, 	arXiv:1903.05265 [cs.IT].


\bibitem{mds1} Kishan Chand Gupta, Sumit Kumar Pandey, Indranil Ghosh Ray, Susanta Samanta. Cryptographically significant mds matrices over finite fields: A brief survey and some generalized results. Advances in Mathematics of Communications, 2019, 13 (4) : 779-843



\bibitem{crp1} Vaudenay S. (1995) On the need for multipermutations: Cryptanalysis of MD4 and SAFER. In: Preneel B. (eds) Fast Software Encryption. FSE 1994. Lecture Notes in Computer Science, vol 1008. Springer, Berlin, Heidelberg.

\bibitem{crp2} H. M. Heys and S. E. Tavares, The design of substitution-permutation networks resistant to differential and linear cryptanalysis, Proceedings of 2nd ACM Conference on Computer and Communications Security, Fairfax, Virginia, 1994, 148{(155)}.
	
\bibitem{crp3} H. M. Heys and S. E. Tavares, Avalanche characteristics of substitution-permutation encryp-	tion networks, IEEE Trans. Comp., 44 (1995), 1131(1139).
		
\bibitem{crp4} H. M. Heys and S. E. Tavares, The design of product ciphers resistant to differential and linear cryptanalysis, Journal of Cryptology, 9 (1996), 1(19).

\bibitem{crp5}V. Rijmen, J. Daemen, B. Preneel, A. Bosselaers and E. D. Win, The cipher SHARK, In 3rd Fast Software Encryption Workshop, LNCS, 1039 (1996), 99(111), Springer-Verlag.

\bibitem{crp6} J. Daemen, L. R. Knudsen and V. Rijmen, The block cipher SQUARE, In 4th Fast Software Encryption Workshop, LNCS, 1267 (1997), 149(165), Springer-Verlag.
	
\bibitem{crp7} J. Daemen and V. Rijmen, The Design of Rijndael: AES - The Advanced Encryption Standard, Springer-Verlag, 2002.

\bibitem{crp8} B. Schneier, J. Kelsey, D. Whiting, D. Wagner, C. Hall and N. Ferguson, Twosh: A 128-bit block cipher, In the rst AES Candidate Conference. National Institute for Standards and Technology, 1998.

\bibitem{crp9}B. Schneier, J. Kelsey, D. Whiting, D. Wagner, C. Hall and N. Ferguson, The Twofish encryption algorithm, Wiley, 1999.

\bibitem{crp10}D. Watanabe, S. Furuya, H. Yoshida, K. Takaragi and B. Preneel, A new keystream generator MUGI, FSE 2002: Fast Software Encryption, Springer Berlin/Heidelberg, 2365 (2002), 179(194).


\bibitem{crp11} G. D. Filho, P. Barreto and V. Rijmen, The Maelstrom-0 Hash Function, In Proceedings of the 6th Brazilian Symposium on Information and Computer Systems Security, 2006.

\bibitem{crp12} P. Gauravaram, L. R. Knudsen, K. Matusiewicz, F. Mendel, C. Rechberger, M. Schlaffer and S. Thomsen, Grostl a SHA-3 Candidate, Submission to NIST, 2008, Available at http:
//www.groestl.info/.

\bibitem{crp13} J. Guo, T. Peyrin and A. Poschmann, The PHOTON family of lightweight hash functions, In CRYPTO, Springer, 2011 (2011), 222(239).

\bibitem{crp14}J. Guo, T. Peyrin, A. Poshmann and M. J. B. Robshaw, The LED block cipher, In CHES	2011, LNCS, 6917 (2011), 32 (341), Springer.


\bibitem{nc1}Hu, Y., Zhang, X., Lee, P.P.C., Zhou, P.
Generalized Optimal Storage Scaling via Network Coding
(2018) IEEE International Symposium on Information Theory - Proceedings, 2018-June, art. no. 8437684, pp. 956-960. 



\bibitem{nc3}Guang, X., Yeung, R.W.
Linear Network Error Correction Coding Revisited
(2020) IEEE International Symposium on Information Theory - Proceedings, 2020-June, art. no. 9174493, pp. 1635-1640. 


\bibitem{nc4}Wu, X., Li, Q., Leung, V.C.M., Ching, P.C.
Joint Fronthaul Multicast and Cooperative Beamforming for Cache-Enabled Cloud-Based Small Cell Networks: An MDS Codes-Aided Approach
(2019) IEEE Transactions on Wireless Communications, 18 (10), art. no. 8786923, pp. 4970-4982. 

\bibitem{nc5}Ko, D., Hong, B., Choi, W.
Probabilistic Caching Based on Maximum Distance Separable Code in a User-Centric Clustered Cache-Aided Wireless Network
(2019) IEEE Transactions on Wireless Communications, 18 (3), art. no. 8638790, pp. 1792-1804. 


\bibitem{nc6}Gu, S., Li, J., Wang, Y., Wang, N., Zhang, Q.
DR-MDS: An Energy-Efficient Coding Scheme in D2D Distributed Storage Network for the Internet of Things
(2019) IEEE Access, 7, art. no. 8648336, pp. 24179-24191. 

\bibitem{nc7}Pedersen, J., Graell Amat, A.I., Andriyanova, I., Brannstrom, F.
Optimizing MDS Coded Caching in Wireless Networks with Device-to-Device Communication
(2019) IEEE Transactions on Wireless Communications, 18 (1), art. no. 8551275, pp. 286-295. 





\bibitem{nc9}Mousavi, S., Zhou, T., Tian, C.
Delayed Parity Generation in MDS Storage Codes
(2018) IEEE International Symposium on Information Theory - Proceedings, 2018-June, art. no. 8437700, pp. 1889-1893. 


\bibitem{nc10}Heidarpour, A.R., Ardakani, M., Tellambura, C.
Network coded cooperation based on relay selection with imperfect CSI
(2018) IEEE Vehicular Technology Conference, 2017-September, pp. 1-5. 


\bibitem{nc11}Heidarpour, A.R., Ardakani, M.
Diversity analysis of MIMO network coded cooperation systems with relay selection
(2018) IEEE Vehicular Technology Conference, 2017-September, pp. 1-6. 


\bibitem{nc12}Liao, J., Wong, K.-K., Zhang, Y., Zheng, Z., Yang, K.
Coding, Multicast, and Cooperation for Cache-Enabled Heterogeneous Small Cell Networks
(2017) IEEE Transactions on Wireless Communications, 16 (10), art. no. 8002600, pp. 6838-6853. 


\bibitem{nc13}Zewail, A.A., Yener, A.
Coded caching for combination networks with cache-aided relays
(2017) IEEE International Symposium on Information Theory - Proceedings, art. no. 8006966, pp. 2433-2437. 


\bibitem{nc14}Guang, X., Fu, F.-W., Zhang, Z.
Variable-Rate Linear Network Error Correction MDS Codes
(2016) IEEE Transactions on Information Theory, 62 (6), art. no. 7447748, pp. 3147-3164. 


\bibitem{nc15}Samadi-Khaftari, V., Esmaeili, M., Gulliver, T.A.
Some Connections Between Classical Coding and Network Coding Over Erroneous Cyclic Networks
(2016) IEEE Access, 4, pp. 5889-5895. 


\bibitem{nc16}Fragouli, C., Soljanin, E.
(Secure) Linear network coding multicast: A theoretical minimum and some open problems
(2016) Designs, Codes, and Cryptography, 78 (1), pp. 269-310. 


\bibitem{nc17}Etzion, T., Storme, L.
Galois geometries and coding theory
(2016) Designs, Codes, and Cryptography, 78 (1), pp. 311-350. 


\bibitem{nc18}Vu, T.X., Duhamel, P., Di Renzo, M.
On the diversity of network-coded cooperation with decode-and-forward relay selection
(2015) IEEE Transactions on Wireless Communications, 14 (8), art. no. 7080912, pp. 4369-4378. 


\bibitem{nc19}Dau, S.H., Song, W., Yuen, C.
On Simple Multiple Access Networks
(2015) IEEE Journal on Selected Areas in Communications, 33 (2), art. no. 6991520, pp. 236-249. 


\bibitem{nc20}Tamo, I., Wang, Z., Bruck, J.
Access versus bandwidth in codes for storage
(2014) IEEE Transactions on Information Theory, 60 (4), art. no. 6737213, pp. 2028-2037. 


\bibitem{nc21}Tamo, I., Wang, Z., Bruck, J.
Zigzag codes: MDS array codes with optimal rebuilding
(2013) IEEE Transactions on Information Theory, 59 (3), art. no. 6352912, pp. 1597-1616. 


\bibitem{nc22}Guang, X., Fu, F.-W., Zhang, Z.
Construction of network error correction codes in packet networks
(2013) IEEE Transactions on Information Theory, 59 (2), art. no. 6320693, pp. 1030-1047. 


\bibitem{nc23}Dai, M., Kwan, H.Y., Sung, C.W.
Linear network coding strategies for the multiple access relay channel with packet erasures
(2013) IEEE Transactions on Wireless Communications, 12 (1), art. no. 6378491, pp. 218-227. 


\bibitem{nc24}Kong, Z., Yeh, E.M., Soljanin, E.
Coding improves the throughput-delay tradeoff in mobile wireless networks
(2012) IEEE Transactions on Information Theory, 58 (11), art. no. 6239596, pp. 6894-6906. 


\bibitem{nc25}Shah, N.B., Rashmi, K.V., Kumar, P.V., Ramchandran, K.
Interference alignment in regenerating codes for distributed storage: Necessity and code constructions
(2012) IEEE Transactions on Information Theory, 58 (4), art. no. 6096412, pp. 2134-2158. 




\bibitem{nc27}Hu, Y., Xu, Y., Wang, X., Zhan, C., Li, P.
Cooperative recovery of distributed storage systems from multiple losses with network coding
(2010) IEEE Journal on Selected Areas in Communications, 28 (2), art. no. 5402494, pp. 268-276. 




\bibitem{nc29}Silva, D., Kschischang, F.R.
Security for wiretap networks via rank-metric codes
(2008) IEEE International Symposium on Information Theory - Proceedings, art. no. 4594971, pp. 176-180. 

\bibitem{nc30}Fragouli, C., Soljanin, E.
Information flow decomposition for network coding
(2006) IEEE Transactions on Information Theory, 52 (3), pp. 829-848. 

\bibitem{crp40} K. C. Gupta and I. G. Ray, Cryptographically significant MDS matrices based on circulant and circulant-like matrices for lightweight applications, Cryptography and Communications, 7 (2015), 257(287).

\bibitem{crp41} M. Liu and S. M. Sim, Lightweight MDS generalized circulant matrices, International Conference on Fast Software Encryption, Lecture Notes in Computer Science, 9783 (2016), 101(120).	Springer, Berlin, Heidelberg.
	
\bibitem{crp42}S. Sarkar AND H. Syed, Lightweight diffusion layer: Importance of Toeplitz matrices, IACR	Trans. Symmetric Cryptol., 2016 (2016), 95(113).
		
\bibitem{crp43}S. Sarkar and H. Syed, Analysis of toeplitz MDS matrices, ACISP 2017, LNCS, 10343 (2017), 3(18).
	


\end{thebibliography}

\documentclass{article}
\usepackage{graphicx}
\usepackage{cite}
\usepackage[left=3cm, right=3cm]{geometry}
\usepackage[latin1]{inputenc}

\usepackage{color}
\usepackage{amsthm,amsmath}
\usepackage{amssymb}
\usepackage{amsfonts}

\newtheorem{theorem}{Theorem}
\newtheorem{definition}{Definition}
\newtheorem{corollary}{Corollary}

\newtheorem{lemma}{Lemma}
\newtheorem{example}{Example}
\newtheorem{remark}{Remark}

\author{Elif Segah Oztas,\\
	\small{Department of Mathematics,}\\
	\small{ Karamanoglu Mehmetbey University, Turkey}\\  \small{esoztas@kmu.edu.tr}}

\begin{document}
\title{Lifted MDS Codes over Finite Fields}
\maketitle
\begin{abstract}

MDS codes are elegant constructions in coding theory and have mode important applications in cryptography, network coding, distributed data storage, communication systems et. In this study, a method is given which MDS codes are lifted to a higher finite field. The presented method satisfies the protection of the distance and creating the MDS code over the $F_q$ by using MDS code over $F_p.$\\
\textit{Keywords:} Lifted MDS codes, distance preserving\\

\end{abstract}
\section{Introduction}

Maximum Distance Separable (MDS) codes \cite{bb} are used across a wide area of modern information technology, cryptography, network coding, data storage etc. In \cite{crp1,crp2,crp3,crp4,crp5,crp6,crp7,crp8,crp9,crp10,crp11,crp12,crp13,crp14},  authors studied on cryptographic approaches. In \cite{nc1,nc3,nc4,nc5,nc6,nc7,nc9,nc10,nc11,nc12,nc13,nc14,nc15,nc16,nc17,nc18,nc19,nc20,nc21,nc22,nc23,nc24,nc25,nc27,nc29,nc30}, MDS codes are studied with network coding.

The main generation method for MDS code is Reed Solomon (RS) codes, especially Generalized Reed Solomon (GRS) codes. In GRS, the code $[n,k,n-k+1]_q$ can obtain where $n \le q$.  
There are some approaches for constructing MDS matrices such that Vandermonde matrix, circulant matrix, Cauchy matrix, Toeplitz matrices etc. \cite{mds1,crp7,crp40,crp41,crp42,crp43,ted}. All of them compute and improve their method over the defined field in the papers. However, calculation complexity increase over the field which has high cardinality for any construction methods for MSD codes, especially in the recursive generating method.

A linear code $C$ with parameters $[n, k, d]_q$ of length $n$ over the finite field $F_q$ where $p$ is a prime and $q$ is a prime power. Any two vectors in $C$ differ in at least $d$
places. Singleton bound is $d \le n-k+1$ and a code satisfying the equality of this bound is called a maximum distance separable (MDS) code.
In this paper, we focus on the extension of existed codes over $F_p$ then we didn't add the MDS code generation method, here. Background on coding theory and related material made be found in \cite{bb}.
In this paper, we give a method to construct MDS codes over $F_q$ ($q=p^t$) by using lift the MDS codes over $F_q$. Moreover, computational complexity is less than other recursive algorithms, and a diversity of the codes are satisfied. These situations give advantages for applications of MDS codes, especially in cryptography

\section{Construction of MDS code over $F_{p^t}$}

In this section, MDS code over $F_p$ ($p$ is a prime) are used to generate MDS codes over $F_{p^t}$ by using distance holder matrix.

\begin{definition}
	Let $M$ is a $n\times n$  diagonal matrix.The entry in the i-th row and j-th column of a matrix $M$ denoted as $m_{ij}$.  $l$ is maximum number of same entries among diagonal entries as follow;
	$$l(M)=\max|\{a : a \in m_{ii} \}|.$$
\end{definition}
For example; $A=\left(\begin{array}{ccc}
	w & 0 &0\\
0 & w^2 &0\\
0 & 0& 1
	
\end{array} \right)
$ and $B=\left(\begin{array}{ccc}
	w & 0 &0\\
	0 & w &0\\
	0 & 0 & 1
\end{array} \right)
$ 
over $F_4$. $l(A)=1$ and $l(B)=2$. In matrix $A$ all diagonal entry has a unique element. In matrix $B$, there are two $w$ and one $1$. Then maximum number of repeated entry in diagonal entries is 2.

\begin{definition}
	Let $M$ be $n\times n$  diagonal matrix  $m_{i,i}\in F_{p^t}^*$ where   $F_{p^t}^*=F_{p^t}-\{0\}$. if $l(M)=1$, it called distance holder matrix. If $l(M)=s$, $M$ is a $s$-distance holder matrix ($s-dh$ matrix).
\end{definition}

The aim of the distance holder matrix is to satisfy the diversity of generators. This means it helps de define different generators for applications of MDS codes. These variations of matrices can be used by the security and multi-node communications systems.

 \begin{lemma}\label{lemmaD}
Let $D$ be an MDS code generation matrix over $F_{p^r}$, then $D'$, obtained by multiplying a row (or column) of $D$ by any element of $F_{p^r}^{*}$, $D$ will also be an MDS code generator matrix.
\end{lemma}

Lemma \ref{lemmaD} is generalized as follows.

\begin{corollary}\label{coro1}
	Let $D$ be an MDS (code generation) matrix, then for any nonsingular diagonal matrices $M_1$ and $M_2$, $M_1 D M_2$ will also be a MDS  matrix.
\end{corollary}

$"\cdot_{F_{p^t}}"$ denote that matrix product operations is over $F_{p^t}$,. 

\begin{definition}
	Let $G$ be a generator matrix of  $C$ that is $[n,k,d]$ $MDS$ code over $F_p$. 
$G\cdot_{F_{p^t}} M$ is generator of $C'$ that is a Lifted MDS code of $C$ over $F_p^t$  
where $M$ is a $n \times n$ $dh$ matrix over $F_{p^t}$ and  $p^t > n$. 
\end{definition}

In the following theorem, codes are lifted to an upper field under some restriction. Then, new codes protect the distance and they are still MDS code over the upper field.

\begin{theorem}\label{teoM}
	Let $C$ be a $[n,k,d]_p$ MDS code over $F_p$. Lifted MDS code of $C$  is $C'$ has parameter $[n,k,d]_{q}$.
\end{theorem}

\begin{proof}
	 Let $C$ be a $[n,k,d]$ MDS code over $F_p$. $d$ has been changed by changing column entries of the generator matrix $G$. In column case, changing the distance is a connected characteristic of $F_p$. Because $p$ is prime, there is no polynomial identification for elements. Then there is no restriction for $d$ except characteristic. Then, operation in field extension to   $F_{p^t}$ that same characteristic as $F_p$ and $p^t > n$  is protect the distance at least $d$.
\end{proof}

By Lemma \ref{lemmaD} and Corollary \ref{coro1}, in same field, the matrix and element operation preserve MDS property. In Theorem \ref{teoM} we satisfy this preservation to the upper finite field by the characteristic of the field.

\begin{remark}{\label{remk1}}
	Diversity of generator matrix are satisfied by using finite field $F_{p^t}$ ($p^t > n$) in Theorem \ref{teoM}. Then number of different generator matrices is $\binom{p^t-1}{n}$. 
\end{remark}

\begin{example}{\label{ex1}}
	Let $G$ be a generator matrix of code $C$  over $F_7$.
	$$G=\left[\begin{array}{cccccccc}
1& 0& 0& 6 &4 &2 &5 &3\\
0 &1 &0 &3 &1 &5 &1 &3\\
0 &0 &1 &3 &5 &2 &4 &6
\end{array}\right]$$
	$C$ is a $[8,3,6]$ MDS code.
	
	Let $M$ be a matrix $dh$-matrix $M=diag(w^{244}, w^{28}, w^{326}, w^{294}, w^{239}, w^{76}, w^{212}, w^{84} )$ over $F_{7^3}.$
	$$G'=G\cdot_{F_{p^t}} M=$$$$=
	\left[\begin{array}{cccccccc}
	    1 &    0  &   0 &w^{221} &w^{223}& w^{288}& w^{253}& w^{239}\\
	    0     &1    & 0 &w^{323} &w^{211} &w^{333} &w^{184} &w^{113}\\
	    0   &  0  &   1 & w^{25} &w^{198} &w^{206}  &   2 &w^{271}	
	\end{array}\right] $$
$G'$ generate a $[8,3,6]$ MDS code over $F_{7^3}$. 

Another example for same code over $F_7$:

Let $M$ be a matrix $dh$-matrix $M=diag( w^{108}, w^{191}, w^{261}, w^{312}, w^{95}, w^{249}, w^{278}, w^{47}  )$ over $F_{7^3}.$

	$$G''=G\cdot_{F_{p^t}} M=$$$$=
\left[\begin{array}{cccccccc}
   1   &  0 &    0&  w^{33}& w^{215} &w^{255}& w^{113}& w^{338}\\
    0  &   1 &    0& w^{178}& w^{246}  &   w&  w^{87}& w^{255}\\
    0  &   0  &   1& w^{108}& w^{119}& w^{102}& w^{245}& w^{299}\\
\end{array}\right] $$
$G''$ generate a $[8,3,6]$ MDS code over $F_{7^3}$.
\end{example}

By Example \ref{ex1}, Distance has been preserved. Moreover, diversity for components of the codes and a MDS code over $F_{7^3} $ have been obtained. By Remark \ref{remk1}, lots of different codes that have the same distance can be generated. This situation has importance in security and communication systems.

\section{Conclusion}

In this paper, we give a method called lifted MDS codes. It satisfies that protection the distance, variation of code components, keep the MDS property in higher finite fields. Moreover, complexity for calculation is less than the previous recursive method which is clear, because there are only matrix multiplications for the generation of new MDS code.

\end{document}